# Decoding Safety Feedback from Diverse Raters: A Data-driven Lens on Responsiveness to Severity

### v4

@@ 4.2 Contrasting with traditional metrics
The way we define and quantify responsiveness addresses these limitations, while offering simplicity and the ability to get insights into how raters utilize different scores on the Likert scale. Nevertheless, as defined in our data model, different factors may influence how safety violations are perceived, regardless of the true underlying severity. While no metric can be an absolute measure of responsiveness to severity, our metrics provide a robust and meaningful measure for evaluating and comparing responsiveness to severity. In appendix C, we conduct extensive simulations to compare our proposed metrics against traditional metrics under different scoring patterns.
@@ 5 Evaluating and comparing responsiveness
To demonstrate the practical utility of our proposed analysis framework with the responsiveness metrics, we apply it to a publicly available pluralistic dataset of safety feedback on AI-generated content. Such safety alignment datasets with granular pluralistic feedback are not readily available.
5.1 Dataset description
We use the dataset of Rastogi et al. (2024), which contains safety feedback from diverse crowd raters for AI-generated images. Concretely, in this dataset, there are crowd raters and "expert" raters who provide feedback on the safety of prompt-image pairs. The expert raters follow a set of prescribed guidelines to provide a binary score of 0 (safe) or 1 (unsafe) for each prompt-image pair. They correspond to trained raters in our framework. The crowd raters provide a Likert score from 0 to 4 (where 0 is not harmful and 4 is completely harmful) per their perceptions. They were recruited based on their demographics. Rastogi et al. identified three demographic axes - gender, ethnicity and age. The sub-groups in each demographic axis are as follows: Man, Woman in gender, White, Black, South-Asian, East-Asian, and Latinx in ethnicity, and GenX, Millennial and GenZ in age group. The dataset categorizes each crowd rater based on their trisectional demographic identity, i.e. their ethnicity, age, and gender. Rastogi et al. distinguish top-level demographic groups, e.g., East-Asian, from trisectional demographic groups, e.g. Black–GenZ–Man. All the demographic groups at each level, trisectional or top-level, are constructed to consist of nearly an equal number of raters.
The dataset contains 5 expert ratings for each prompt-image pair; on average, 4.09 expert raters give the same rating per prompt-image pair. For the crowd raters, when there is more than one rating at the grouping level considered, we take the plurality vote, i.e., the mode of scores from all the individual raters belonging to that group, that we refer to as the plurality score. Taking the plurality vote preserves the ordinal nature of the scale. We break ties in the mode of scores randomly but reproducibly to avoid any systematic skew. The distribution of scores provided by the crowd raters is shown in appendix D along with group cohesion metrics for each of the top-level demographic groups. Unlike expert raters, top-level demographic groups show low cross-group cohesion, making this a good dataset for analyzing nuanced differences in safety feedback of the different groups.
(a) Ethnicity
(b) Age
(c) Gender
5.2 Results by trisectional demographic groups
@@
Figure 2: Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, and auroc for trisectional demographic groups of crowd raters when binary reference $U$ is obtained from expert raters. All confidence intervals are within ±$\pm 0.01$. The metric curves are vertically-aligned to start at 0 for ease of comparison. Legend gives the values by which the curves are translated.
@@
Figure 2 shows monotonic precision area (MPA), weighted recall area (WRA), their harmonic mean (HM), Kendall’s $\tau$, and AUROC for trisectional demographic groups of crowd raters when binary reference $U$ is obtained from expert raters, i.e., guideline-based reference. We note that the trisectional groups comprising Latinx and East-Asian ethnicities, aka Latinx trisections and East-Asian trisections, consistently have the lowest MPA ($p$-value < 0.05 under permutation test), and consequently, the lowest HMs as well. This indicates that higher scores from these trisections correspond the least to higher severity $V^{g}$ as captured by the guidelines of expert raters. Figure 1(a) further validates the same as we note for instance that the East-Asian top-level demographic group does not exhibit consistent gains in $Precision(S)$ when plurality score $S$ goes from 1 to 3. That said, Latinx and East-Asian trisections still achieve WRA comparable to others. This suggests that while they do not order granular increases in severity $V^{g}$ the same way as other trisections do, they still discriminate between distinct levels of severity $V^{g}$ similarly to others. We note that Kendall’s $\tau$ and AUROC track WRA closely as both of them capture aspects of concordance. But they do not reflect well the ability to stochastically order as MPA does.
Figure 3 shows metrics for the trisectional demographic groups when the binary reference $U$ is obtained from crowd raters themselves, excluding the group being evaluated. We binarize the scores of crowd raters using boundaries $k\in[1,4]$ and report the metrics aggregated over all the boundaries via micro-averaging to ensure robustness. Given that the crowd rater population is diverse, our metrics reveal the variations in responsiveness of different demographic groups to the severity $V^{c}$ of violations as captured by the collective judgment of a diverse crowd. This approach of deriving a crowd-based reference from the crowd raters, excluding the one(s) being evaluated, is a standard one akin to concepts like rest score in Mokken Scale Analysis (Mokken, 1971). Here, we also provide metrics from Item Response Theory (IRT) and Mokken Scale Analysis (MSA) with the unidimensional latent trait being the true severity of prompt-image pairs. They are now applicable because now we are analyzing ordinal scale ratings relative to a criterion derived from those ratings.
@@
Figure 3: Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, auroc, Mokken h, and irt discrimination $\alpha$ for trisectional demographic groups of crowd raters when binary reference $U$ is obtained from crowd raters, excluding the group being evaluated. All confidence intervals are within ±$\pm 0.01$. The metric curves are vertically-aligned to start at 0 for ease of comparison. Legend gives the values by which the curves are translated.
@@
MPA, WRA, and the traditional metrics continue to exhibit the same behaviors as before. Again, IRT and MSA only provide insights into raters’ ability to discriminate between levels of the latent trait based on the discrimination $\alpha$ or H values, but not the ability to stochastically order. Unlike before, we note that now the trisectional groups comprising Black ethnicity, aka Black trisections, consistently have the lowest MPA and also the lowest WRA ($p$-values < 0.05). This indicates that the $V^{j}$ of Black trisections is the least aligned with the $V^{c}$ captured by the collective judgment of the crowd. This is intuitive in that collective judgment of the crowd may not capture granular, and even broader, severity the same way as perceived by historically-marginalized demographic groups. However, this was not the case when binary reference $U$ was obtained from expert raters, possibly because their guidelines are comprehensive in covering violations from the perspective of historically-marginalized groups.
5.3 Results by violation types
We now look at the trends for the entire crowd rater population on the three different violation types in prompt-image pairs, namely, bias, sexual, and violent, when binary reference $U$ is obtained from expert raters. Figure 4 presents $Precision(S)$ and $Recall(S)$ curves for the three violation types and figure 4 gives MPA, WRA, their harmonic mean (HM), Kendall’s $\tau$, and AUROC for the three violation types. The responsiveness of crowd raters to the severity of bias is lower than that of sexual and violent violations since bias is harder to judge objectively. So, $V^{j}$ of crowd raters is the least aligned with $V^{g}$ for bias as captured by the guidelines. Appendix E further provides metrics for top-level demographic groups on violation types, and appendix F contains some qualitative examples.
Figure 4: (a) $Precision(S)$ and $Recall(S)$ at plurality scores $S=0$ to $4$ for the entire crowd rater population on the three violation types when binary reference $U$ is obtained from expert raters.
(b) Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, and auroc for the entire crowd rater population. All confidence intervals are within ±$\pm 0.01$.
6 Recommendations and limitations
@@
Gathering granular feedback from diverse raters belonging to various different sociocultural groups is a very expensive and meticulous task, often more so than simply gathering feedback via prescribed guidelines (Rastogi et al., 2024; Kirk et al., 2024). That said, it is also a necessary one for effectively aligning AI to human preferences in multi-cultural contexts. Granular safety feedback gathered from diverse raters can be used to fulfill two different objectives during the alignment process:
  1. 1.
@@
    To validate and improve guideline-driven safety alignment. Here, our metrics can highlight whenever there is granular and/or broader differences between $V^{j}$ of a group (top-level, trisectional, or otherwise) and $V^{g}$, i.e., when MPA and/or WRA are low. Furthermore, our metrics can contribute to improving the quality of feedback during the fine-tuning or RL optimization stages through a 2-step action plan: identify raters who exhibit the highest responsiveness to severity from each and every group (top-level, trisectional, or otherwise), and curate optimization datasets with items at varying levels of severity based on the scores of these raters (Bergman et al., 2024).
  2. 2.
    To directly perform safety alignment. Here, our findings show that naively optimizing for the collective judgment of the crowd may not effectively reflect the perceptions of all the groups. Firstly, our metrics can identify the groups whose granular or broader perceptions of severity may not be captured well in the collective judgment of a diverse crowd, i.e., low MPA or low WRA or both, be it collective judgment that is aggregated naively or otherwise (e.g., social choice). Further, our metrics can identify groups who show similar responsiveness to severity, which can potentially reduce the need to gather feedback from all such groups, and hence, reduce the costs.
We note that our metrics are descriptive and non-parametric in nature. So, while they enable a deeper interpretation of nuanced differences in observed trends, they cannot directly estimate the causal parameters behind the trends. Such estimations would require bigger datasets and more complex structural modeling, both of which we leave to future works.
@@ 7 Conclusion
We formulated non-parametric metrics to analyze the nuanced differences in safety feedback of raters expressed via ordinal scales, addressing the limitations in existing approaches. Applying these metrics to a study involving diverse crowd raters, we found significant variations in how different demographic groups respond to the severity of violations when providing safety feedback on AI-generated content. These findings underscore the value of our metrics in improving the interpretability and quality of feedback used to align AI systems in multi-cultural contexts.
@@ A.1 Proof for stochastic ordering property
given that $S$ and $T$ are independent by construction. Since $P(T=t)$ is non-negative and $P(V>T|S=s_{1},T=t)\geq P(V>T|S=s_{2},T=t)$ for every $t\in\mathcal{T}$, hence
@@ A.2 Proof for discrimination property
Let $a(t)=P(S\geq s,U=1,T=t)$, $b(t)=P(U=1,T=t)$, and $c(t)=P(T=t)$. Since $S$ and $T$ are independent by construction, $P(S\geq s,T=t)=P(S\geq s)c(t)$. Then,
@@
Appendix B Metric implementation
@@
{python}
import numpy, sklearn
def get_mpa_wra_for_rater( likert_scores: list, binary_reference: list, likert_length: int ) -> tuple[float]: """ Args: - likert_scores: array of Likert scores from a crowd rater - binary_reference: array of binary reference U - likert_length: length of the Likert scale [0, …, K]
Returns - tuple of floats for the crowd rater (MPA, WRA, harmonic mean) """ # convert to numpy arrays reference_arr = numpy.array(binary_reference) scores_arr = numpy.array(likert_scores)
one_counts, zero_counts, counts = [], [], [] for k in range(0, likert_length): scores_flag = scores_arr == k count_at_k = scores_flag.sum() one_count_at_k = (reference_arr * scores_flag).sum() zero_count_at_k = count_at_k - one_count_at_k counts.append(count_at_k) one_counts.append(one_count_at_k) zero_counts.append(zero_count_at_k)
# compute Y_so at Likert scores and then MPA y_so = [] norm = numpy.ceil(likert_length / 2) * numpy.floor(likert_length / 2) for k in range(0, likert_length): demr, numr = numpy.array(counts[:k]), numpy.array(one_counts[:k]) prev_precisions = numpy.maximum.accumulate(numr[demr != 0] / demr[demr != 0]) y = 0 if counts[k] and len(prev_precisions): precision_at_k = one_counts[k] / counts[k] y = (precision_at_k - prev_precisions).sum() y_so.append(y) y_so.append(0.) mpa = sklearn.metrics.auc(range(likert_length + 1), y_so) / norm
# compute Y_d at Likert scores and then WRA y_d = [] U_0, U_1 = (1 - reference_arr).sum(), reference_arr.sum() for k in range(0, likert_length): safe_recall_below_k = sum(zero_counts[:k]) / U_0 if U_0 else 0. unsafe_recall_at_k = one_counts[k] / U_1 if U_1 else 0. y = safe_recall_below_k * unsafe_recall_at_k y_d.append(y) y_d.append(0.) wra = sklearn.metrics.auc(range(likert_length + 1), y_d)
hm = (2 * mpa * wra) / (mpa + wra) return (mpa, wra, hm)
Appendix C Comparison of metrics via simulations

### v5

@@ 4.3 Contrasting with other metrics
Looking at the parametric alternatives, we note that IRT models, e.g., Graded Response Model, impose the assumption of monotonicity with respect to some common latent scale(s), obscuring the non-monotonic empirical trends. In other words, the diverse perceptions $V^{j}$ of raters are all projected onto a single scale $V$, reducing pluralistic variations to misfit or residual error. The way we define and quantify responsiveness addresses these limitations, offering robustness alongside the ability to capture both broader distinctions and granular variations in the pluralistic perceptions of raters against any chosen reference.
@@ 5 Evaluating and comparing responsiveness
To demonstrate the practical utility of our proposed analysis framework with the responsiveness metrics, we apply it to two very different publicly available pluralistic datasets of safety feedback. Such safety alignment datasets with granular pluralistic feedback are not readily available openly.
5.1 Dataset descriptions
Dataset 1. We experiment with the dataset of Rastogi et al. (2025; 2024), which contains safety feedback from diverse crowd raters for AI-generated images. Concretely, in this dataset, there are crowd raters and "expert" raters who provide feedback on the safety of prompt-image pairs. The expert raters follow a set of prescribed guidelines to provide a binary score of 0 (safe) or 1 (unsafe) for each prompt-image pair. They correspond to trained raters in our framework. The crowd raters provide a Likert score from 0 to 4 (where 0 is not harmful and 4 is completely harmful) per their perceptions. They were recruited based on their demographics. Rastogi et al. identified three demographic axes, gender, ethnicity and age. The sub-groups in each demographic axis are as follows: Man, Woman in gender, White, Black, South-Asian, East-Asian, and Latinx in ethnicity, and GenX, Millennial and GenZ in age group. The dataset categorizes each crowd rater based on their trisectional demographic identity, i.e., their ethnicity, age, and gender. Rastogi et al. distinguish top-level demographic groups, e.g., East-Asian, from trisectional demographic groups, e.g. Black–GenZ–Man. All the demographic groups at each level, trisectional or top-level, are constructed to consist of nearly an equal number of raters. Further details of the dataset are available in appendix E.
Dataset 2. We further experiment with the toxicity dataset of Kumar et al. (2021) that is different in scale, domain, modality, and demographic axes. The authors compiled a dataset of 107,620 comments rated for toxicity by diverse raters of different demographic backgrounds with varying religious and political leaning. Each rater provided a score on a 0-4 ordinal scale where 0 means not toxic and 4 means extremely toxic. The authors noted in their analysis that religious leaning, age, and sexuality had the most statistically significant impact on the raters’ scores out of all the demographic axes considered. Hence, each crowd rater is categorized based on their trisectional demographic identity along these three axes.
5.2 Results
@@
Results on Dataset 1. Figure 1 shows monotonic precision area (MPA), weighted recall area (WRA), their harmonic mean (HM), Kendall’s $\tau$, and AUROC for trisectional demographic groups of crowd raters when binary reference $U$ is obtained from expert raters, i.e., guideline-based reference. We note that the trisectional groups comprising Latinx and East-Asian ethnicities, referred to as Latinx trisections and East-Asian trisections, consistently have the lowest MPA ($p$-value < 0.05 under permutation test), and consequently, the lowest HMs as well. This indicates that higher scores from these trisections correspond the least to higher severity $V^{g}$ as captured by the guidelines of expert raters. Figure 8(a) in appendix E further validates the same as we note, for instance, that the East-Asian top-level demographic group does not exhibit consistent gains in $Precision(s)$ when plurality score $S$ goes from 1 to 3. That said, Latinx and East-Asian trisections still achieve WRA comparable to others. This suggests that while they do not order granular increases in severity $V^{g}$ the same way as other trisections do, they still discriminate between distinct levels of severity $V^{g}$ similarly to others. We note that Kendall’s $\tau$ and AUROC track WRA closely (Pearson $r>0.7$) as both of them capture aspects of concordance. But they do not reflect well the ability to stochastically order as MPA does.
@@
Figure 1: Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, and auroc for trisectional demographic groups of crowd raters when binary reference $U$ is obtained from expert raters. All confidence intervals are within $\pm 0.01$. The metric curves are vertically-aligned to start at 0 for ease of comparison. Legend gives the values by which the curves are translated.
Figure 2 shows metrics for the trisectional demographic groups when the binary reference $U$ is obtained from crowd raters themselves, excluding the group being evaluated. We obtain the binary reference $U$ using boundaries $\{1,2,3,4\}$ and report the metrics macro-averaged over all the boundaries to ensure robustness. Trends remain the same with other aggregations. Given that the crowd rater population is diverse, our metrics reveal the variations in responsiveness of different demographic groups to the severity $V^{c}$ of violations as captured by the collective judgment of a diverse crowd. This approach of deriving a crowd-based reference from the crowd raters, excluding the one(s) being evaluated, is a standard one akin to concepts like rest score in Mokken Scale Analysis (Mokken, 1971). Here, we also provide metrics from IRT (Graded Response Model) and Mokken Scale Analysis (MSA) with the latent trait being the true severity $V$ of prompt-image pairs. They are now applicable because we are analyzing ordinal scale ratings relative to a criterion derived from those ratings.
Figure 2: Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, auroc, Mokken h, and irt discrimination $\alpha$ for trisectional demographic groups of crowd raters when binary reference $U$ is obtained from crowd raters, excluding the group being evaluated. All confidence intervals are within $\pm 0.01$. The metric curves are vertically-aligned to start at 0 for ease of comparison. Legend gives the values by which the curves are translated.
@@
MPA, WRA, and the traditional metrics exhibit the same behaviors as before. IRT and MSA reflect how broader distinctions in the perceptions of raters align with broader distinctions in the assumed latent trait via the $\alpha$ and H values, but not granular variations. Unlike before, we note that trisectional groups comprising Black ethnicity, referred to as Black trisections, now consistently have the lowest MPA and also the lowest WRA ($p$-values < 0.05). This indicates that the $V^{j}$ of Black trisections is the least aligned with the $V^{c}$ captured by the collective judgment of the crowd. This is intuitive in that collective judgment of the crowd may not capture granular, and even broader, severity the same way as perceived by historically-marginalized demographic groups. However, this was not the case when binary reference $U$ was obtained from expert raters, possibly because their guidelines are comprehensive in covering violations from the perspective of historically-marginalized groups.
Lastly, in appendix F, we provide metrics for top-level demographic groups per the three violation types in the dataset (bias, sexual explicitness, violence) and present some qualitative examples.
Results on Dataset 2. In figure 3, we show monotonic precision area (MPA), weighted recall area (WRA), the harmonic mean (HM), Kendall’s $\tau$, AUROC, and IRT (Graded Response Model) for trisectional groups of raters based on their religious leaning, sexuality, and age group. As before, we obtain the binary reference $U$ from the crowd raters, excluding the rater group being evaluated, using boundaries in $\{1,2,3,4\}$ and report the metrics macro-averaged over the boundaries to ensure robustness. Once again, we note that Kendall’s $\tau$, AUROC, and IRT discrimination $\alpha$ continue to track WRA closely (Pearson $r>0.7$) as all of them capture discrimination. But again, they do not reflect well the ability to stochastically order as MPA does (Pearson $r<0.1$).
Figure 3: Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, auroc, and irt discrimination $\alpha$ for trisectional groups of diverse raters. All confidence intervals are within $\pm 0.01$. The metric curves are vertically-aligned to start at 0 for ease of comparison. Legend gives the values by which the curves are translated.
Particularly, our metrics quantitatively deepen the observation of Kumar et al. (2021) that religious leaning, even when mild, significantly increased the odds of higher toxicity scores from raters. We observe that trisections where religion is important consistently have lower MPA ($p$-value < 0.05 under permutation test), and consequently, lower HMs as well. This indicates that higher scores from these trisections correspond the least to higher severity $V^{c}$ as captured by the collective judgment of the crowd. The same trends hold for trisections with LGBTQ+ as sexuality, where our metrics again quantitatively elucidate the same observation of Kumar et al. Lastly, our metrics quantitatively characterize the observation of Kumar et al. (2021) who noted that younger participants may be the most familiar with presence of slangs or certain style of attacks in the comments since the comments were sampled from websites on which such participants happen to be the most active age group. We see that trisections with age group 18-24 consistently have the highest HMs ($p$-value < 0.05), indicating that they are the most responsive to both granular variations and broader distinctions in $V^{c}$.
6 Alignment in pluralistic settings
@@
Gathering nuanced feedback from diverse raters belonging to various different sociocultural groups is a very expensive and meticulous task, often more so than simply gathering feedback via prescribed guidelines (Wang et al., 2025; Rastogi et al., 2024; Kirk et al., 2024). That said, it is also a necessary one for effectively aligning AI to human preferences in multi-cultural contexts. Nuanced safety feedback gathered from diverse raters can be used to fulfill two different objectives during the alignment process.
@@
To validate and improve guideline-driven safety alignment. Here, our metrics can highlight whenever there is granular and/or broader differences between $V^{j}$ of a group (trisectional, top-level, or otherwise) and $V^{g}$, i.e., when MPA and/or WRA are low. Furthermore, our metrics can contribute to improving the quality of feedback during the fine-tuning or RL optimization stages through a 2-step action plan: select raters who exhibit the highest responsiveness to severity from each and every group (trisectional, top-level, or otherwise), and curate optimization datasets with items at varying levels of severity based on the scores of these raters (Bergman et al., 2024).
To directly perform safety alignment. Before training a reward model (RM) on the gathered data, we need to understand the granular and/or broader differences in the feedback of the groups (trisectional, top-level, or otherwise) because our findings show that naively optimizing for the collective judgment of the crowd may not effectively capture the preferences of all the groups. For groups that exhibit low MPA and low WRA when the reference $U$ is obtained from crowd raters, their $V^{j}$ is misaligned with $V^{c}$. In such cases, we inspect the items they rated high vs. low severity in a pairwise manner to determine if those reflect some unique broader perceptions of safety or not. Then, such items can accordingly be included or excluded during the optimization of the RM, be it pairwise optimization or pointwise. Alternatively, some groups may have high WRA but low MPA. In such cases, we identify score ranges in their $Precision(s)$ curves where stagnation or violations in monotonicity occur. In these ranges, granular increases in $V^{j}$ do not indicate increases in $V^{c}$. We inspect the items at scores in these ranges to verify that they indeed reflect unique granular perceptions, then include them with appropriate weights in the optimization of the RM to reinforce diverse granular preferences.
@@ 7 Conclusion
We formulated non-parametric metrics to analyze the nuanced differences in safety feedback of raters expressed via ordinal scales, addressing the limitations in existing approaches. Applying these metrics to case studies involving diverse crowd raters, we found significant variations in how different demographic groups respond to the severity of violations when providing safety feedback. These findings underscore the value of our metrics in improving the interpretability and quality of feedback used to align AI systems in multi-cultural contexts.
Broader Impact Statement
Descriptive non-parametric approaches like ours as well as model-based approaches like Item Response Theory cannot explain why exogenous differences exist in the real world between groups. This is a fundamental limitation of all non-qualitative approaches. Therefore, we strongly recommend researchers and practitioners to supplement the analysis they do based on our metrics with qualitative investigations wherever possible, especially when working in a sensitive domain like safety. Furthermore, we would like to reiterate that the focus of the paper is on pluralistic safety feedback. Therefore, our metrics are not designed as a means to exclude any groups who show lower responsiveness to a chosen reference. In fact, we strongly encourage researchers and practitioners to use our metrics as a means to identify when responsiveness varies, and carefully curate more inclusive objectives for optimization.
Lastly, we treat reproducibility as a very important aspect, especially in sensitive areas like safety. In appendix C, we have provided the implementation of the metrics in Python that we used for our experiments. Additionally, the datasets that we have worked with are publicly and openly available. Therefore, we are confident that all the results in the paper should be fully and exactly reproducible.
@@ A.1 Proof for stochastic ordering property
given that $S$ and $T$ are independent. Since $P(T=t)$ is non-negative and $P(V>T|S=s_{1},T=t)\geq P(V>T|S=s_{2},T=t)$ for every $t\in\mathcal{T}$, hence
@@ A.2 Proof for discrimination property
Let $a(t)=P(S\geq s,U=1,T=t)$, $b(t)=P(U=1,T=t)$, and $c(t)=P(T=t)$. Since $S$ and $T$ are independent, $P(S\geq s,T=t)=P(S\geq s)c(t)$. Then,
@@
Appendix B Recommendations and visualizations for observed trends
In table 1, we provide some recommendations on steps to take for specific trends that may be observed in monotonic precision area (MPA) and weighted recall area (WRA) of individual crowd raters or rater groups when the binary reference $U$ is either obtained from trained raters (capturing $V^{g}$) or from crowd raters themselves (capturing $V^{c}$).
Possible reason(s) Recommended steps
MPA is low Not all scores of the ordinal scale are used by the rater or the rater group due to the presence of some response bias. $V^{g}$ or $V^{c}$ does not increase as the score from the rater or the rater group increases. Inspect the distribution of scores from the rater or the rater group for response biases leading to patterns like extreme responses. Encourage the rater or the rater group to utilize the full scale. Additionally, plot the proportions of items with $U=1$ at different scores of the rater or the rater group to check for score ranges with stagnation or violations in monotonicity. Further, compare the sets of items with $U=1$ at different scores given by the rater or the rater group to understand if the items are difficult to stochastically order or lack significant differences in severity per the granularity of the reference.
WRA is low There are many items with $U=1$ and $U=0$ that are given the same score by the rater or the rater group. Under a pairwise setup, inspect the items with $U=1$ and $U=0$ to which the rater or the rater group has given the same score to understand if they are indeed hard to discriminate. Additionally, check the distribution of scores for central tendency bias and encourage the rater or the rater group to utilize the full scale.
Appendix C Metric implementation
@@
def get_mpa_wra_for_rater(
likert_scores: list,
binary_reference: list,
likert_length: int
) -> tuple[float]:
Args:
- likert_scores: array of Likert scores from a crowd rater
- binary_reference: array of binary reference U
- likert_length: length of the Likert scale [0, ..., K]
Returns
- tuple of floats for the crowd rater (MPA, WRA, harmonic mean)
# convert to numpy arrays
reference_arr = numpy.array(binary_reference)
scores_arr = numpy.array(likert_scores)
one_counts, zero_counts, counts = [], [], []
for k in range(0, likert_length):
scores_flag = scores_arr == k
count_at_k = scores_flag.sum()
one_count_at_k = (reference_arr * scores_flag).sum()
zero_count_at_k = count_at_k - one_count_at_k
counts.append(count_at_k)
one_counts.append(one_count_at_k)
zero_counts.append(zero_count_at_k)
# compute Y_so at Likert scores and then MPA
y_so = []
norm = numpy.ceil(likert_length / 2) * numpy.floor(likert_length / 2)
for k in range(0, likert_length):
demr, numr = numpy.array(counts[:k]), numpy.array(one_counts[:k])
prev_precisions = numpy.maximum.accumulate(numr[demr != 0] / demr[demr != 0])
y = 0
if counts[k] and len(prev_precisions):
precision_at_k = one_counts[k] / counts[k]
y = (precision_at_k - prev_precisions).sum()
y_so.append(y)
y_so.append(0.)
mpa = sklearn.metrics.auc(range(likert_length + 1), y_so) / norm
# compute Y_d at Likert scores and then WRA
y_d = []
U_0, U_1 = (1 - reference_arr).sum(), reference_arr.sum()
for k in range(0, likert_length):
zero_recall_below_k = sum(zero_counts[:k]) / U_0 if U_0 else 0.
one_recall_at_k = one_counts[k] / U_1 if U_1 else 0.
y = zero_recall_below_k * one_recall_at_k
y_d.append(y)
y_d.append(0.)
wra = sklearn.metrics.auc(range(likert_length + 1), y_d)
hm = (2 * mpa * wra) / (mpa + wra)
return (mpa, wra, hm)
Appendix D Comparison of metrics via simulations